OpenStudy (anonymous):

313. The total was $29.00 after adding a $4.00 tip to the bill for a family of five, each of whom ordered a hot dog. What was the price of a hot dog? (1 point) 4 – 5x = 29; $5.00 4 + 5x = 29; $5.00 29 – 5x = 4; $5.00 5x = 29 = 4; $5.00

OpenStudy (texaschic101):

4 + 5x = 29 5x = 29 - 4 5x = 25 x = 5